# Symmetric Preparation of Systems

### v2

@@ 2. Formal power series solutions
where $u$ and $m$ are analytic and $m=m^{*}$, which gives analytic symmetric $A_{j}$, j=0, 1j=0,\,1. We shall show that this differential is surjective in a neighborhood of $(0,0)$. We find from (3.4) that $m=U_{0}^{-1}A_{0}(U_{0}^{-1})^{*}$ and
$2t\operatorname{Re}U(t,x)u^{*}(t,x)=\int_{0}^{t}A_{1}(s,x)\,ds+R(x)-U(t,x)m(x)U^{*}(t,x)$
@@
for some analytic symmetric $R(x)$. Since the right hand side has to vanish at $t=0$, we take $R(x)=U_{0}(x)m(x)U_{0}^{*}(x)=A_{0}(x)$ and then
$2\operatorname{Re}U(t,x)u^{*}(t,x)=\frac{1}{t}\left(\int_{0}^{t}A_{1}(s,x)\,ds+A_{0}(x)-U(t,x)m(x)U^{*}(t,x)\right)=A(t,x)$
@@
which is analytic and symmetric. By substituting $m$ in $A$ we find that $A$ depends linearly on $A_{j}$, j=0, 1j=0,\,1, and there exists an analytic and antisymmetric $B(t,x)$ so that $2Uu^{*}=A+B$, which gives the solution
(3.5) $u(t,x)=\frac{1}{2}(A(t,x)-B(t,x))(U^{-1})^{*}(t,x)$
@@
Thus (3.4) is surjective so by the inverse function theorem we obtain a local right inverse to $\mathcal{F}$ and local surjectivity of the nonlinear mapping (3.1) in a neighborhood $V$ of $(U,0)$. This neighborhood depend on $U(t,x)$ and thus by $F(t,x)$ by (3.3). Observe that $\mathcal{F}$ will also be surjective on some $(F_{1},F_{0})$ such that $F_{0}(0)\neq 0$.
@@
For any symmetric $M(x)$ such that $M(0)=0$, there exists a neighborhood $\omega$ of $(0,0)$ such that $(U,M)\in V$ at $\omega$, and then (3.5) gives a solution to (3.1) in $\omega$. But (3.4) is not injective, since different antisymmetric $B$ gives different solutions by (3.5). Actually, we have for fixed $U$ and $M\equiv 0$ that the mapping $(A_{0},A_{1},B)\mapsto(u,m)$ is bijective near $(0,0)$.
@@
It remains to prove uniqueness in the case $U(t,x)$ is symmetric and positive definite. Since $m$ already is determined by (3.4), we only need to prove that there is a unique symmetric solution $u(t,x)$ to equation (3.5). In that case, we obtain from Lemma A.1 a unique analytic antisymmetric $B(t,x)$ that solves
@@
The smooth $N\times N$ systems form a local Fréchet spaces with the $C^{\infty}(\overline{\omega})$ norms, which are tame if $\overline{\omega}$ is a compact manifold with boundary, see Theorem C.8. For tame Fréchet spaces one can use the Nash–Moser theorem, which is the inverse function theorem for tame Fréchet spaces, see Theorem C.12. The difference to the analytic case is that is it not enought to check the local surjectivity of the linearized map when $M\equiv 0$, instead one has to check that for any $U$ and $M$ in a neighborhood.
@@
As before, it suffices to show that if $|U(0,0)|\neq 0$, $M(x)$ is symmetric and $M(0)=0$ then the following $C^{\infty}$ mapping is surjective near $(0,0)$
(4.1) $C^{\infty}\ni(U(t,x),M(x))\overset{\mathcal{F}}{\longrightarrow}\\ \left(\partial_{t}\big(U(t,x)(t\operatorname{Id}_{N}+M(x))U^{*}(t,x)\big),U_{0}(x)M(x)U_{0}^{*}(x)\right)\\ =(F_{1}(t,x),F_{0}(x))\in C^{\infty}$
@@
where $U_{0}(x)=U({0,x})$. Here $F_{0}$ and $F_{1}$ are symmetric, $F_{1}(0,0)>0$ and $F_{0}(0)=0$. By (3.2) we then obtain $F$ from $F_{0}$ and $F_{1}$. It follows from Example C.4 that the mapping $\mathcal{F}$ is tame of degree 1.
@@
For any such $(F_{1},F_{0})$ we may take smooth $(U(t,x),M(x))$ so that $U(0,0)U^{*}(0,0)=F_{1}(0,0)>0$ and $M(0)=0$. Since we only need local solutions, we may assume that $U(t,x)\in C^{\infty}(\overline{\omega}_{n+1})$ so that $U(0,0)$ is invertible, $M(x)\in C^{\infty}(\overline{\omega}_{n})$ is symmetric and $M(0)=0$, where $\overline{\omega}_{k}\subset\mathbf{R}^{k}$ is a compact neighborhood of the origin.
@@
Let $V$ be a bounded neighborhood of $(U(t,x),M(x))$ in $C^{\infty}(\overline{\omega}_{n+1})\times C^{\infty}(\overline{\omega}_{n})$ so that the first component is invertible and the second component is symmetric. Since $M(0)=0$ we may for any $\varepsilon>0$ choose $V$ and a compact neighborhood $\overline{\omega}$ of the origin so that $(\overline{U}(t,x),\overline{M}(x))\in V$ gives $\|\overline{M}(x)\|\leq\varepsilon$ when $x\in\overline{\omega}$. For any $(U(t,x),M(x))\in V$ the differential of $\mathcal{F}$ at $(U(t,x),M(x))$ is equal to
(4.2) $C^{\infty}\ni(u(t,x),m(x))\overset{d\mathcal{F}}{\longrightarrow}\\ \left(\partial_{t}\big(2\operatorname{Re}(U(t\operatorname{Id}_{N}+M(x))u^{*})+UmU^{*}\big),U_{0}mU_{0}^{*}+2\operatorname{Re}(U_{0}Mu_{0}^{*})\right)\\ =(A_{1}(t,x),A_{0}(x))\in C^{\infty}$
@@
where $u_{0}(x)=u(0,x)$, and $m$, $A_{0}$ and $A_{1}$ are symmetric. We shall show that if $\overline{\omega}_{n+1}$ and $\overline{\omega}_{n}$ are sufficiently small, then the mapping $d\mathcal{F}$ is tame and surjective with a tame right inverse, which will show the surjectivity of $\mathcal{F}$ by Theorem C.12. For that, we need that $\|M(x)\|\ll 1$ for $|x|\ll 1$.
We find from (4.2) that
(4.3) $m=U_{0}^{-1}\Big(A_{0}-2\operatorname{Re}(U_{0}Mu^{*}_{0})\Big)(U_{0}^{-1})^{*}$
is linear in $Mu_{0}^{*}$ and if $P(t,M(x))=t\operatorname{Id}_{N}+M(x)$ then
(4.4) $2\operatorname{Re}UP(t,M(x))u^{*}=\int_{0}^{t}A_{1}(s,x)\,ds-U(t,x)m(x)U^{*}(t,x)+R(x)=A(t,x)$
@@
for some smooth symmetric $R(x)$. Observe that $A$ depends linearly on $m$ and thus linearly on $Mu_{0}^{*}$ by (4.3). By (4.4) there exists a smooth antisymmetric $B(t,x)$ so that
(4.5) $2P(t,M(x))u^{*}(t,x)=\\ U^{-1}(t,x)\int_{0}^{t}A_{1}(s,x)\,ds-m(x)U^{*}(t,x)+U^{-1}(t,x)R(x)+U^{-1}(t,x)B(t,x)$
@@
Since $R(x)$ is symmetric and $B(0,x)$ is antisymmetric, we may replace $R(x)$ with $R(x)+B(0,x)$ and let $B(0,x)\equiv 0$. Then we can replace $B(t,x)$ with $tB(t,x)$, which gives
(4.6) $U^{-1}(t,x)tB(t,x)=P(t,M(x))U^{-1}(t,x)B(t,x)-M(x)U^{-1}(t,x)B(t,x)$
Now we shall use the symmetric Malgrange division theorem given by Corollary B.4 on every term on the right hand side of (4.5) using (4.6). That will give a $C^{\infty}$ surjective tame map with a $C^{\infty}$ tame right inverse by Example C.13.
First, if $P=P(t,M(x))$ then Corollary B.4 gives smooth $C_{1}(t,x)$ and $R_{1}(x)$ so that
(4.7) $U^{-1}(t,x)\int_{0}^{t}A_{1}(s,x)\,ds=PC_{1}(t,x)+R_{1}(x)\qquad\text{near $(0,0)$}$
where $C_{1}$ and $R_{1}$ depend linearly on $m$ and thus linearly on $Mu_{0}^{*}$ by (4.3), and that
(4.8) $m(x)U^{*}(t,x)=PC_{2}(t,x)+R_{2}(x)\qquad\text{near $(0,0)$}$
where $C_{2}$ and $R_{2}$ are smooth and depend linearly on $Mu_{0}^{*}$. We also find
(4.9) $U^{-1}(t,x)=PC_{3}(t,x)+R_{3}(x)\qquad\text{near $(0,0)$}$
@@
with smooth $C_{3}$ and $R_{3}$ such that $R_{3}(0)=U^{-1}(0,0)$ is invertible, thus $R_{3}(x)$ is invertible when $|x|\ll 1$. Finally, we obtain smooth $C_{4}(t,x)$ and $R_{4}(x)$ so that
(4.10) $M(x)U^{-1}(t,x)B(t,x)=PC_{4}(t,x)+R_{4}(x)\qquad\text{near $(0,0)$}$
and $C_{4}(t,0)=R_{4}(0)=0$. Here $C_{4}$ and $R_{4}$ depend linearly on $B$ but the differentials with respect to $B$ is $\mathcal{O}(M)$. We have that $C_{j}$ and $R_{j}$ depend tamely on $U$ and $M$, $\forall\,j$.
Summing up, we find from (4.6)–(4.10) that
(4.11) $2Pu^{*}(t,x)=P(C_{1}+C_{2}+C_{3}R+C_{4}+U^{-1}B)+R_{1}+R_{2}+R_{3}R+R_{4}=PC_{0}+R_{0}$
@@
near $(0,0)$, where $C_{0}(t,x)$ and $R_{0}(x)$ are smooth depending linearly on $R$ $Mu_{0}^{*}$ and $B$. Thus for fixed $U$, $M$ and $B$, we can write
(4.12) $\left\{\begin{aligned} &C_{0}=C_{00}+C_{01}Mu_{0}^{*}+C_{02}R+C_{03}B+U^{-1}B\\ &R_{0}=R_{00}+R_{01}Mu_{0}^{*}+R_{3}R+R_{03}B\end{aligned}\right$.
@@
where $R_{3}$ and $U^{-1}$ are invertible. Observe that, except $U^{-1}B$ and $R_{3}R$, the products are not matrix multiplications, instead they represent linear maps. We have that the norms $\|C_{03}\|$ and $\|R_{03}\|$ are bounded by $C\|M\|$.
@@
Now, since $u_{0}(x)$ appears in the equation we shall let $G(t,x)=G_{0}(x)+tG_{1}(t,x)$ for any smooth $G$, and $C_{jk}(t,x)=C_{jk}^{0}(x)+tC_{jk}^{1}(t,x)$, $\forall\,jk$. For fixed $U$, $M$ and $B$ we obtain that $2u^{*}\equiv C_{0}$ with $R_{0}\equiv 0$ if
(4.13) $\left\{\begin{aligned} &2u_{0}^{*}-C_{01}^{0}Mu_{0}^{*}-C_{02}^{0}R\equiv C_{00}^{0}+(C_{03}B)_{0}+(U^{-1}B)_{0}\\ &-tC_{01}^{1}Mu_{0}^{*}+2tu_{1}^{*}-tC_{02}^{1}R\equiv tC_{00}^{1}+t(C_{03}B)_{1}+t(U^{-1}B)_{1}\\ &R_{01}Mu_{0}^{*}+R_{3}R\equiv-R_{00}-R_{03}B\end{aligned}\right$.
By dividing the second equation by $t$, this system has the following matrix for $(u^{*}_{0},u_{1}^{*},R)$
(4.14) $\begin{pmatrix}2\operatorname{Id}_{N}-C_{01}^{0}M&0&-C_{02}^{0}\\ -C_{01}^{1}M&2\operatorname{Id}_{N}&-C_{02}^{1}\\ R_{01}M&0&R_{3}\end{pmatrix}$
which is invertible when $M\equiv 0$ since $R_{3}$ is invertible, and therefore when $\|M(x)\|\ll 1$. Thus, if the neighborhood $V$ is small enough, then for any $B$ we obtain a unique solution $u^{*}\equiv u_{0}^{*}+tu_{1}^{*}\equiv C_{0}/2$ with $R_{0}\equiv 0$ in a sufficiently small compact neighborhood $\overline{\omega}$ of $(0,0)$. This gives a smooth tame map by Example C.4, and the neighborhood $\overline{\omega}$ only depends on the bounds on $U$, $M$ and $B$.

### v3

@@ 2. Formal power series solutions
We shall show that this differential is surjective for such systems in a neighborhood of $(0,0)$. We find from (3.3) that ${\bf m}=T_{m}(U^{-1}A_{0}(U^{-1})^{*})$ and
$2t^{m}\operatorname{Re}U(t,x){\bf u}^{*}(t,x)=I_{m}(A_{1})(t,x)+\mathbb{R}(t,x)-U{\bf m}U^{*}(t,x)$
@@
for some analytic symmetric $\mathbb{R}(t,x)$ such that $\partial_{t}^{m}\mathbb{R}\equiv 0$. Since the right hand side has to vanish of order $m$ at $t=0$, we find that $\mathbb{R}(t,x)=T_{m}(U{\bf m}U^{*})(t,x)=A_{0}(t,x)$ and then
$2\operatorname{Re}U(t,x){\bf u}^{*}(t,x)=t^{-m}\Big(I_{m}(A_{1})(t,x)+A_{0}(t,x)-U{\bf m}U^{*}(t,x)\Big)=A(t,x)$
@@
which is analytic and symmetric. By substituting $\bf m$} in $A$ we find that $A$ depends linearly on $A_{j}$, j=0, 1j=0,\,1, and that there exists an analytic and antisymmetric $B(t,x)$ so that $2U{\bf u}^{*}=A+B$, which gives $u^{*}=U^{-1}(A+B)/2$ and the solution
(3.4) ${\bf u}(t,x)=\frac{1}{2}(A(t,x)-B(t,x))(U^{-1})^{*}(t,x)$
@@
This gives the surjectivity of $d\mathcal{F}$ in (3.3) and by the inverse function theorem we obtain local surjectivity of the nonlinear mapping $\mathcal{F}$ in (3.1) in a neighborhood $V$ of $(U,0)$. Observe that $\mathcal{F}$ will also be surjective on some $(F_{1},F_{0})$ such that $|F_{0}(t,0)|\ll 1$.
@@
For any symmetric ${\mathbb{M}}(t,x)$ such that $\partial_{t}^{m}{\mathbb{M}}(t,x)\equiv 0$ and ${\mathbb{M}}(t,0)=0$, there exists a neighborhood $\omega$ of $(0,0)$ such that $(U,{\mathbb{M}})\in V$ at $\omega$, and then (3.4) gives a solution to (3.1) in $\omega$. But (3.3) is not injective, since any antisymmetric $B$ gives a different solution. Actually, we have for fixed $U$ and ${\mathbb{M}}\equiv 0$ that the mapping $(A_{0},A_{1},B)\mapsto({\bf u},{\bf m})$ is bijective near $(0,0)$.
@@
It remains to prove uniqueness in the case $U(t,x)$ is positive definite near $(0,0)$, which we obtain when $U(0,0)=(F_{1}(0,0)/m!)^{1/2}>0$. Since $\bf m$} already is determined, we shall prove that there is only one symmetric solution ${\bf u}(t,x)$ to equation (3.4). If $\operatorname{Im}u^{*}=0$ then we obtain from Lemma A.1 a unique analytic antisymmetric $B(t,x)$ that solves
@@
The smooth $N\times N$ systems form a local Fréchet spaces with the $C^{\infty}(\overline{\omega})$ norms, which are tame if $\overline{\omega}$ is a compact neighborhood with boundary, see Theorem B.8. For tame Fréchet spaces one can use the Nash–Moser theorem, which is the inverse function theorem for tame Fréchet spaces. The difference to the analytic case is that is it not enought to check the local surjectivity of the linearized map when ${\mathbb{M}}(0)\equiv 0$, instead one has to check that for any smooth $U$ and $\mathbb{M}$} in a neighborhood, see Theorem B.12.
@@
So we shall show that if $U(t,x)\in C^{\infty}$, $|U(0,0)|\neq 0$, ${\mathbb{M}}(t,x)=\sum_{j=0}^{m-1}t^{j}M_{j}(x)\in C^{\infty}$ is symmetric such that $M_{j}(0)=0,\ \forall\,j$, then the following $C^{\infty}$ mapping is surjective near $(0,0)$
(4.1) $C^{\infty}\ni(U(t,x),{\mathbb{M}}(t,x))\overset{\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}^{m}\big(U(t,x)(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x))U^{*}(t,x)\big),T_{m}(U{\mathbb{M}}U^{*})(t,x))\Big)\\ =(F_{1}(t,x),F_{0}(t,x))\in C^{\infty}$
@@
Here $F_{j}$ is symmetric, j=0, 1,j=0,\,1, $F_{1}(0,0)>0$, $\partial_{t}^{m}F_{0}\equiv 0$ and $F_{0}(t,0)\equiv 0$. As before, if we obtain a solution $(F_{1},F_{0})$ to (4.1) then we obtain $F(t,x)$ from (3.2) so that $F_{0}(t,x)=T_{m}(F)(t,x)$ and $F_{1}(t,x)=\partial_{t}^{m}F(t,x)$. It follows from Example B.4 that the mapping $\mathcal{F}$ is tame of degree $m$.
@@
Now, for any such $(F_{1}(t,x),F_{0}(t,x))$ we may choose $(U_{0}(t,x),{\mathbb{M}}_{0}(t,x))$ so that $U_{0}(t,x)$ is invertible and $U_{0}(0,0)U^{*}_{0}(0,0)\equiv F_{1}(0,0)$ and ${\mathbb{M}}_{0}(t,x)$ is a symmetric polynomial in $t$ of degree less than $m$ such that ${\mathbb{M}}_{0}(t,0)\equiv 0$. Observe that we may choose symmetric $U_{0}$ such that $U_{0}(0,0)=F_{1}^{1/2}(0,0)>0$, and then $U_{0}>0$ near the origin. Since we shall only need local solutions, we may assume that $U_{0}(t,x)$ and ${\mathbb{M}}_{0}(t,x)\in C^{\infty}(\overline{\omega})$ where $\overline{\omega}$ is compact neighborhood of the origin so that $U(t,x)$ is invertible and ${\mathbb{M}}(t,x)$ is symmetric such that $\partial_{t}^{m}{\mathbb{M}}(t,x)\equiv 0$ and ${\mathbb{M}}(t,0)\equiv 0$.
@@
Then we let $V$ be a bounded neighborhood of $(U_{0}(t,x),{\mathbb{M}}_{0}(t,x))$ in $C^{\infty}(\overline{\omega})$ so that the first component is invertible and the second component is symmetric polynomial in $t$ of degree less than $m$. If $U_{0}>0$ then $V$ is a neighborhood with symmetric matrices. Since ${\mathbb{M}}_{0}(t,0)\equiv 0$ we may for $\varepsilon>0$ choose $V$ and a compact neighborhood $\overline{\omega}$ so that $(U,{\mathbb{M}})\in V$ gives $\|{\mathbb{M}}\|\leq\varepsilon$ in $\overline{\omega}$.
We find for any $(U(t,x),{\mathbb{M}}(t,x))\in V$ that the differential of $\mathcal{F}$ at $(U(t,x),{\mathbb{M}}(t,x))$ is equal to
(4.2) $C^{\infty}\ni({\bf u}(t,x),{\bf m}(t,x))\overset{d\mathcal{F}}{\longrightarrow}\\ \Big(\partial_{t}^{m}\big(2\operatorname{Re}(U(t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t,x)){\bf u}^{*})+U{\bf m}U^{*}\big),T_{m}(U{\bf m}U^{*})+2\operatorname{Re}T_{m}(U{\mathbb{M}}{\bf u}^{*})\Big)\\ =(A_{1}(t,x),A_{0}(t,x))\in C^{\infty}$
where ${\bf u}(t,x)$ and ${\bf m}(t,x)$ are smooth such that $\bf m$} is symmetric and $\partial_{t}^{m}{\bf m}\equiv 0$ which gives smooth symmetric $A_{j}$, j=0, 1j=0,\,1, such that $\partial_{t}^{m}A_{0}\equiv 0$.
@@
Next, we shall consider the special case when the functions are constant in $x$. We shall show that if ${\bf u}(t)$ and ${\bf m}(t)\in C^{\infty}(\overline{\omega})$ for a sufficiently small neighborhoods $\overline{\omega}$ of the origin in $\mathbf{R}$, then the mapping $d\mathcal{F}$ is tame and surjective with a tame right inverse. For that, the condition that $\|{\mathbb{M}}\|\ll 1$ in $\overline{\omega}$ is necessary.
Proposition 4.1.
Let $U(t)$ and ${\mathbb{M}}(t)$ be smooth and $N\times N$ valued functions such that $|U(t)|\neq 0$, $\partial^{m}_{t}{\mathbb{M}}(t)\equiv 0$ and $\|{\mathbb{M}}(t)\|\ll 1$. Then in a sufficiently small compact neighborhood $\overline{\omega}$ of the origin, the mapping $d\mathcal{F}$ given by (4.2) is surjective, having a tame local right inverse of degree 3 in $\overline{\omega}$. If $U(t)$ is positive definite, then $d\mathcal{F}$ is bijective and has a tame local inverse of degree 3.
By Theorem B.12, we obtain that the $C^{\infty}$ tame mapping $\mathcal{F}$ is surjective on a neighborhood $W$ of a given $(F_{1}(t),F_{0}(t))$ close to the origin, having a smooth tame right inverse $\mathcal{F}^{-1}$ giving $(U(t),{\mathbb{M}}(t))=\mathcal{F}^{-1}(F_{1}(t),F_{0}(t))$. Now, if $F_{j}(t)=F_{j}(t,0)$ then for $|x|\ll 1$ we obtain that $(F_{1}(t,x),F_{0}(t,x))\in W$, so we obtain the symmetric preparation (1.7) with
(4.3) $(U(t,x),{\mathbb{M}}(t,x))=\mathcal{F}^{-1}(F_{1}(t,x),F_{0}(t,x))$
in a neighborhood of $(0,0)$.
If $U(t)$ is positive definite, then by Theorem B.12 we obtain that the mapping $\mathcal{F}$ is a bijection in a neighborhood $W$ of a given $(F_{1}(t),F_{0}(t))$ near the origin with the tame inverse $\mathcal{F}^{-1}$ giving a unique $(U(t),{\mathbb{M}}(t))=\mathcal{F}^{-1}(F_{1}(t),F_{0}(t))$. If $F_{j}(t)=F_{j}(t,0)$ and $|x|\ll 1$ then we obtain from (4.3) the unique symmetric solution $(U(t,x),{\mathbb{M}}(t,x))$ to (1.7) near the origin which finishes the proof of Theorem 1.1 in the smooth case. ∎
It remains to prove Proposition 4.1. We find from (4.2) that
(4.4) ${\bf m}=T_{m}\left(U^{-1}\big(A_{0}-2\operatorname{Re}(U{\mathbb{M}}{\bf u}^{*})\big)(U^{-1})^{*}\right)$
and if $P(t,{\mathbb{M}}(t))=t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t)$ with symmetric ${\mathbb{M}}(t)=\sum_{j<m}t^{j}M_{j}$ we have
(4.5) $2\operatorname{Re}UP(t,{\mathbb{M}}(t)){\bf u}^{*}=I_{m}(A_{1})(t)-U(t){\bf m}(t)U^{*}(t)+\mathbb{R}(t)=A(t)$
@@
for some smooth symmetric $\mathbb{R}(t)$ such that $\partial_{t}^{m}\mathbb{R}(t)\equiv 0$. Observe that $A$ depends linearly on $\bf m$} by (4.4) and thus linearly on $T_{m}({\mathbb{M}}{\bf u}^{*})$, and that we may assume that $\|{\mathbb{M}}\|=\sum_{j<m}\|M_{j}\|\ll 1$. By (4.5) there exists a smooth antisymmetric $B(t)$ so that
(4.6) $2P(t,{\mathbb{M}}(t)){\bf u}^{*}(t)=U^{-1}(t)I_{m}(A_{1})(t)-{\bf m}(t)U^{*}(t)+U^{-1}(t)\mathbb{R}(t)+U^{-1}(t)B(t)$
@@
We shall keep $B$ fixed in the following, and we may take $B(t)\equiv 0$. Since $\mathbb{R}(t)$ is symmetric and $B(t)$ is antisymmetric, we may replace $\mathbb{R}(t)$ with $\mathbb{R}(t)-T_{m}(B)(t)$ and let $T_{m}(B)\equiv 0$. Then we can replace $B(t)$ by $t^{m}B(t)$ which gives
(4.7) $U^{-1}(t)t^{m}B(t)=P(t,{\mathbb{M}}(t))U^{-1}(t)B(t)-{\mathbb{M}}(t)U^{-1}(t)B(t)$
We shall use the following result, which is the adjoint version of Proposition 3.2 in [3] in the special case when $\pi_{m}=\operatorname{Id}_{N}$.
Proposition 4.2.
Let $G(t)\in\mathcal{S}(\mathbf{R})$ be $N\times N$ valued and let $P(t,{\mathbb{M}}(t))=t^{m}\operatorname{Id}_{N}+{\mathbb{M}}(t)$ with $N\times N$ system ${\mathbb{M}}(t)=\sum_{j<m}t^{j}M_{j}$, ${\mathbb{M}}=\left\{\,M_{j}\,\right\}_{j<m}$ and $\|{\mathbb{M}}\|=\sum_{j<m}\|M_{j}\|$. Then we can find $N\times N$ systems $Q(t,{\mathbb{M}},G)\in C^{\infty}$ and $\mathbb{R}(t,{\mathbb{M}},G)=\sum_{j<m}t^{j}R_{j}({\mathbb{M}},G)\in C^{\infty}$ depending linearly on $G(t)$, such that
(4.8) $G(t)=P(t,{\mathbb{M}}(t))Q(t,{\mathbb{M}},G)+\mathbb{R}(t,{\mathbb{M}},G){\quad}\text{if}{\quad}\|{\mathbb{M}}\|<1$
Also, we have the following estimates when $\|{\mathbb{M}}\|<1$
(4.9) $\|\partial_{t}^{{\alpha}}\partial_{{\mathbb{M}}}^{{\beta}}Q(t,{\mathbb{M}},G)\|\leq C_{{\alpha}{\beta}}\int(\|G\|+\|G^{(j)}\|)\,dt{\quad}j={\alpha}+3+m(|{\beta}|+1)$
$\|\partial_{{\mathbb{M}}}^{{\beta}}R_{k}(t,{\mathbb{M}},G)\|\leq C_{{\beta}}\int(\|G\|+\|G^{(j)}\|)\,dt{\quad}j=2+m(|{\beta}|+1)\qquad k<m$
We find that (4.8) gives a smooth surjective tame map $C^{\infty}\ni(Q,\mathbb{R})\mapsto G\in\mathcal{S}$ of degree 0 having a smooth tame right inverse of degree $3+m$, see Example B.14. Observe that the estimate (4.8) is not tame in the parameters $\mathbb{M}$}. Thus, for $\|{\mathbb{M}}\|\ll 1$ we shall use Proposition 4.2 on every term on the right hand side of (4.6) using (4.7), after cut off so that they have compact support. We will denote by $\mathbb{R}_{k}(t)$ the terms that are polynomials of degree less than $m$ in $t$.
First, if $P=P(t,{\mathbb{M}}(t))$ then Proposition 4.2 gives smooth $C_{1}(t)$ and $\mathbb{R}_{1}(t)$ so that
(4.10) $U^{-1}(t)I_{m}(A_{1})(t)=PC_{1}(t)+\mathbb{R}_{1}(t)\qquad\text{near $t=0$}$
where $C_{1}$ and $\mathbb{R}_{1}$ are smooth depending linearly on $\bf m$} and thus linearly on $T_{m}({\mathbb{M}}{\bf u}^{*})$ by (4.4), and that
(4.11) ${\bf m}(t)U^{*}(t)=PC_{2}(t)+\mathbb{R}_{2}(t)\qquad\text{near $t=0$}$
where $C_{2}$ and $\mathbb{R}_{2}$ are smooth and depending linearly on $T_{m}({\mathbb{M}}{\bf u}^{*})$. We also find
(4.12) $U^{-1}(t)=PC_{3}(t)+\mathbb{R}_{3}(t)\qquad\text{near $t=0$}$
@@
with smooth $C_{3}$ and $\mathbb{R}_{3}$, thus $U^{-1}\mathbb{R}=PC_{3}\mathbb{R}+\mathbb{R}_{3}\mathbb{R}$. By (4.9), $C_{3}$ is uniformly bounded when $\|{\mathbb{M}}\|<1$. Thus $\mathbb{R}_{3}(0)$ is invertible if $\|{\mathbb{M}}(0)\|\ll 1$ since $\mathbb{R}_{3}(0)=U^{-1}(0)-{\mathbb{M}}(0)C_{3}(0)$. Finally, we obtain smooth $C_{4}(t)$ and $\mathbb{R}_{4}(t)$ so that
[…]
where $\|C_{4}\|_{k}$ and $\|\mathbb{R}_{4}\|_{k}$ are $\mathcal{O}(\|{\mathbb{M}}\|)$ for any local $C^{k}$ norm. We find from Proposition 4.2 that $C_{j}(t)$ and $\mathbb{R}_{j}(t)$, $\forall\,j$, depend smoothly and tamely of degree $3+m$ on $U(t)$ and ${\mathbb{M}}(t)$.
Summing up, we find from (4.7), and (4.10)–(4.13) that
(4.14) $2Pu^{*}(t)=P\left(C_{1}+C_{2}+C_{3}\mathbb{R}+C_{4}B+U^{-1}B\right)+\mathbb{R}_{1}+\mathbb{R}_{2}+\mathbb{R}_{3}\mathbb{R}+\mathbb{R}_{4}B=PC_{0}+\mathbb{R}_{0}$
@@
locally near the origin in $\mathbb{R}$. Here $C_{0}$ and $\mathbb{R}_{0}$ are smooth depending linearly on $\mathbb{R}$, $B$ and $T_{m}({\mathbb{M}}{\bf u}^{*})$. Thus for fixed $U$ and $\mathbb{M}$} we can write
(4.15) $\left\{\begin{aligned} &C_{0}=C_{00}+C_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+C_{3}\mathbb{R}+C_{4}B+U^{-1}B\\ &\mathbb{R}_{0}=\mathbb{R}_{00}+\mathbb{R}_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+\mathbb{R}_{3}\mathbb{R}+\mathbb{R}_{4}B\end{aligned}\right$.
@@
where $\mathbb{R}_{3}(t)$ is invertible near $t=0$ and $\|C_{4}\|_{k}+\|R_{4}\|_{k}=\mathcal{O}(\|{\mathbb{M}}\|)$, $\forall\,k$. Observe that the products are matrix multiplications, except $C_{01}T_{m}({\mathbb{M}}{\bf u}^{*})$ and $R_{01}T_{m}({\mathbb{M}}{\bf u}^{*})$, which represent linear maps.
@@
Now, since $T_{m}({\mathbb{M}}{\bf u}^{*})$ appears in the equations we shall put $G(t)=T_{m}(G)(t)+t^{m}G_{1}(t)$ for $G\in C^{\infty}$ and $C_{jk}(t)=T_{m}(C_{jk})(t)+t^{m}C_{jk}^{1}(t)$, $\forall\,jk$. Then for fixed $U$, $\mathbb{M}$} and $B$ we obtain that $2{\bf u}^{*}\equiv C_{0}$ and $\mathbb{R}_{0}\equiv 0$ if
(4.16) $\left\{\begin{aligned} &2T_{m}({\bf u}^{*})-T_{m}(C_{01}T_{m}({\mathbb{M}}{\bf u}^{*}))-T_{m}(C_{3}\mathbb{R})\equiv T_{m}(C_{00})+T_{m}(C_{4}B)+T_{m}(U^{-1}B)\\ &-C_{01}^{1}T_{m}({\mathbb{M}}{\bf u}^{*})+2{\bf u}_{1}^{*}-(C_{3}\mathbb{R})_{1}\equiv C_{00}^{1}+(C_{4}B)_{1}+(U^{-1}B)_{1}\\ &\mathbb{R}_{01}T_{m}({\mathbb{M}}{\bf u}^{*})+\mathbb{R}_{3}\mathbb{R}\equiv-\mathbb{R}_{00}-\mathbb{R}_{4}B\end{aligned}\right$.
Observe that the first equation is modulo $\mathcal{O}(t^{m})$, and that
(4.17) $\|T_{m}({\mathbb{M}}{\bf u}^{*})\|_{k}\lesssim\|{\mathbb{M}}\|\|{\bf u}^{*}\|_{k+m-1}$
since $\|{\mathbb{M}}\|_{k}\lesssim\|{\mathbb{M}}\|$ and $\|T_{m}{\bf u}^{*}\|_{k}\lesssim\|{\bf u}^{*}\|_{k+m-1}$ for any local $C^{k}$ norm. Then we get from (4.17) that (4.18) is given by
(4.18) $\begin{pmatrix}2\operatorname{Id}_{N}&0&-T_{m}(C_{3}\ \boldsymbol{\cdot})\\ 0&2\operatorname{Id}_{N}&-(C_{3}\ \boldsymbol{\cdot})_{1}\\ 0&0&\mathbb{R}_{3}\end{pmatrix}$
modulo terms having local $C^{k}$ norms that are $\mathcal{O}(\|{\mathbb{M}}\|)$. Here $\operatorname{Im}(C\ \boldsymbol{\cdot})$ is the mapping $E\mapsto\operatorname{Im}(CE)$ and $T_{m}(C\ \boldsymbol{\cdot})$ is the mapping $E\mapsto T_{m}(CE)$. The system (4.18) is invertible close to $t=0$ since $\mathbb{R}_{3}(0)$ is invertible. Thus, we find that the system (4.16) is invertible in a compact neighborhood $\overline{\omega}$ of the origin if $\|{\mathbb{M}}\|\ll 1$, which can be obtained by shrinking $V$ and $\overline{\omega}$. Since this system is linear, smooth and invertible, we find from Example B.4 that it has a tame smooth inverse of degree 0 which gives a solution $(T_{m}({\bf u}^{*}),{\bf u}^{*}_{1},\mathbb{R})$.